OpenStudy (ajanijones):

A number cube with the numbers 1, 2, 3, 4, 5, and 6 is rolled 1000 times. About how many times would it be expected that a number greater than 4 is rolled? A. 500 B. 333 C. 667 D. 167

OpenStudy (anonymous):

lol "number cube" = dice

OpenStudy (anonymous):

numbers greater than 4, 5 and 6 chance per roll of getting 5 = 1/6, of getting 6 = 1/6 of 5 or 6 = 1/6 + 1/6 = 1/3 1000 times = 1000*1/3 = 333

OpenStudy (anonymous):

b